OpenStudy (anonymous):

describes the system of equations 2x - y + 4 = 0 and 4x - 2y + 7 = 0

OpenStudy (shinalcantara):

2x - y +4 = 0 ----equation 1 4x - 2y + 7 = 0 ----equation 2 multiplying 2 to equation 1 you'll have: 2x - y + 4 =0 4x - 2y + 8 = 0 ---equation 1' subtract equation 1' from 2 4x - 2y + 7 = 0 - 4x - 2y + 8 = 0 0x + 0y -1 = 0 -1 is not equal to zero the system of equations doesn't have a solution